OpenStudy (anonymous):

Consider the integral ∫2,0∫sqrt(4−y),0f(x,y)dxdy. If we change the order of integration we obtain the sum of two integrals: ∫ba∫g2(x)g1(x)f(x,y)dydx+∫dc∫g4(x)g3(x)f(x,y)dydx a= b= g1(x)= g2(x)= c= d= g3(x)= g4(x)=

hartnn (hartnn):

can you plot the region ? x= 0 to x = sqrt (4-y) y =0 to y = 2

OpenStudy (anonymous):

i did...trying to get the image. give me a second

hartnn (hartnn):

x= sqrt (4-y) will be the right half of parabola x^2 = 4-y

hartnn (hartnn):

something like this ?